Cost per Equivalent Unit

The following information concerns production in the Forging Department for November. All direct materials are placed into the process at the beginning of production, and conversion costs are incurred evenly throughout the process. The beginning inventory consists of $43,050 of direct materials.

ACCOUNT Work in Process—Forging Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Nov. 1 Bal., 3,500 units, 60% completed 51,030
30 Direct materials, 32,000 units 387,200 438,230
30 Direct labor 52,870 491,100
30 Factory overhead 73,022 ? 564,122
30 Goods finished, ? units ?
30 Bal., 2,800 units, 60% completed ?

a. Determine the number of units transferred to the next department.
units

b. Determine the costs per equivalent unit of direct materials and conversion. If required, round your answer to two decimal places.

Cost per equivalent unit of direct materials $
Cost per equivalent unit of conversion $

c. Determine the cost of units started and completed in November.
$
Answer #1
a
Number of units transferred 32700 =3500+32000-2800
b
Direct materials Conversion
To complete beginning work in process 0 1400 =3500*(1-60%)
Units started and completed 29200 29200 =32700-3500
Ending work in process 2800 1680 =2800*60%
Equivalent units 32000 32280
Direct materials Conversion
Current period costs 387200 125892
Divide by Equivalent units 32000 32280
Cost per equivalent unit 12.10 3.90
Cost per equivalent unit of direct materials 12.10
Cost per equivalent unit of conversion 3.90
c
Units started and completed 29200
X Cost per equivalent unit 16 =12.10+3.90
Cost of units started and completed 467200